At 250C the vapor pressure of mercury is 74.4 torr and at 320C the vapor pressure is 376.3 torr. Find the \Delta Hvap in KJ/mol.
Answer #1

T1 = 25.0 oC

=(25.0 + 273)K

= 298.0 K

T2 = 32.0 oC

=(32.0 + 273)K

= 305.0 K

P1 = 74.4 torr

P2 = 3.763*10^2 torr

use:

ln(P2/P1) = (Hvap/R)*(1/T1 - 1/T2)

ln(3.763*10^2/74.4) = ( Hvap/8.314)*(1/298.0 - 1/305.0)

1.620931 = (Hvap/8.314)*(7.702*10^-5)

Hvap = 174982 J/mol

Hvap = 175 KJ/mol

Answer: 175 KJ/mol